OpenStudy (anonymous):

product of (5x+3)(5x-3)

OpenStudy (anonymous):

What is X?

OpenStudy (anonymous):

It is any number, basically nothing.

OpenStudy (anonymous):

I know what the definition is, but usually they give it during questions.

OpenStudy (anonymous):

Yeah, but I almost never get that, so I just pretend it is nothing.

OpenStudy (lyrae):

Multiplication by conjugate \[(a + b)(a - b) = a^2 - b^2\] Example \[(2+7x)(2-7x) = 2^2 - (7x)^2 = 4 - 49x\]

OpenStudy (lyrae):

Correction \[4 - 49x^2\]

OpenStudy (anonymous):

\[5x*5x= 25x ^{2}\] \[5x*3= 15x\] \[-3*5x= 15x\] \[-3*3= -9\] \[25x ^{2}-15x+15x-9\] \[=25x ^{2}-9\]